# Bootstrapping the critical behavior of multi-matrix models

### v1

@@ Abstract
Given a matrix model, by combining the Schwinger-Dyson equations with positivity constraints on its solutions, in the large $N$ limit one is able to obtain explicit and numerical bounds on its moments. This technique is known as bootstrapping with positivity. In this paper we use this technique to estimate the critical points and exponents of several matrix multi-models. As a proof of concept, we first show it can be used to find the well-studied quartic single matrix model’s critical phenomena. We then apply the method to several similar “unsolved” 2-matrix models with various quartic interactions. We conjecture and present strong evidence for the string susceptibility exponent for some of these models to be $\gamma=1/2$, which heuristically indicates that the continuum limit will likely be the Continuum Random Tree. For the other 2-matrix models, we find estimates of new string susceptibility exponents that may indicate a new continuum limit. We then study an unsolved 3-matrix model that generalizes the 3-colour model with cubic interactions. Additionally, for all of these models, we are able to derive explicitly the first several terms of the free energy in the large $N$ limit as a power series expansion in the coupling constants at zero by exploiting the structure of the Schwinger-Dyson equations.
@@ 1 Introduction
The primary quantities of interest are:
@@
Typically, one says that a model is solved (to leading order) if all these quantities can be written in terms of the coupling constants in the large $N$ limit. In general, these models are notoriously difficult to solve, even to leading order. For single matrix models with polynomial potentials and a small class of multi-matrix potentials, corrections to these models can be solved recursively using a process known as Topological Recursion [eynard2016counting, eynard2007invariants]. For general multi-matrix models, very little is known and much of is lacks a rigorous basis, with the exceptions of works like [guionnet2004first, guionnet2005combinatorial]. This is largely due to a shortage of analytic techniques. For an overview of what is known, see [kazakov2000solvable, eynard2011formal].
Of further interest is the study of the critical phenomena and a potential continuum limit of these models. Often, the quantities of interest have an asymptotic expansion around a critical point, i.e., a configuration of the coupling constants where one of the above quantities fails to be analytic. As a formal series that enumerates maps, such an expansion can be understood in terms of the asymptotic growth of the number of faces in the maps associated with the model. Such asymptotic expansions were first discovered in the ’90s and connected to 2D conformal field theory [di19952d, witten1990two]. This connection comes from studying the statistical properties of planar maps constructed from polygons of a given model as random metric spaces. Such measures are viewed as discretized measures over piecewise flat geometries used to construct models of 2d Euclidean quantum gravity. One then hopes to take appropriate limits so that the size of each polygon tends to zero and that these random metric spaces converge in some sense to a continuum theory. In the past twenty years, much work has been done to rigorously study such continuum limits [angel2003growth, chassaing2004random, le2010scaling]. Analogous ideas are used with Random Tensors [bonzom2012random, lionni2018colored] and Random Dynamical Triangulations [agishtein1992critical, loll2019quantum] for constructing toy models of higher dimensional gravity. It is worth noting that, despite matrix models corresponding to two-dimensional surfaces, multi-matrix models appear in both these frameworks as models with a space-time dimension greater than two [ambjorn2001lorentzian, lionni2018colored].
@@
where $f$ is a polynomial function and $2-\gamma\not\in\mathbb{Z}.$ This expansion is often referred to as the double scaling limit. The constant $\gamma$ is known as the string susceptibility exponent [di19952d]. When $\gamma=-1/2$, the associated continuum limit is expected to be the Brownian map [le2013uniqueness, marckert2006limit]. When $\gamma=1/2$, the associated continuum limit is expected to be the Continuum Random Tree [aldous1991continuumI, aldous1991continuumII, aldous1993continuumIII] which is known as branched polymers in the physics literature [cates1985fractal]. While these two are the most common and well-studied, other exponents are possible [kazakov1989appearance, korchemsky1992matrix]. In particular, this is possible when the model studied has multiple coupling constants that can be fine-tuned [ambjorn2016generalized, distler19902d]. Note that critical exponents do not always imply the continuum limit, but rather give a heuristically strong indication of them. For an example, see [kazakov1986ising].
To combat the above lack of analytical methods to solve multi-matrix models, the method of bootstrapping with positivity was first introduced in the innovative work [Lin2020]. Through this method, one derives stricter and stricter bounds on the moments of matrix models by combining the relations of the Schwinger-Dyson equations with positivity constraints on the spectral measure. Solving for these bounds is a non-linear optimization problem. In particular, in [kazakov2022analytic, zheng2023bootstrap] it was shown to be applicable to multi-matrix models that are unsolved, providing much-needed insight. In the same works, the relaxation bootstrap method was introduced, turning this non-linear problem into a linear one. Even more recently in [li2024analytic], further progress has been made in studying the 2-matrix model from [kazakov2022analytic] by adding an ansatz for the structure of moments. In [perez2024loop], positivity constraints on matrix integrals over the unitary group were used for bootstrapping. Bootstrapping with positivity has also been used in the previous work of the first two authors to study matrix models motivated by Noncommutative Geometry [hessam2022bootstrapping], ultimately leading to the recent analytic solution for many moments and the free energy of these models [khalkhali2024coloured]. The idea of bootstrapping with positivity constraints has also been used in other areas of mathematical physics such as matrix quantum mechanics [aikawa2022bootstrap, berenstein2022bootstrapping, berenstein2023semidefinite, bhattacharya2021numerical, tchoumakov2021bootstrapping], lattice gauge theory [kazakov2024bootstrap, kazakov2023bootstrap], Feynman integrals [zeng2023feynman], the Ising model on the lattice [cho2022bootstrapping], and more.
@@ 1.1 Outline of main results
In this paper, we will apply the bootstrapping with positivity technique to study the critical points and exponents of single, two, and 3-matrix models. In [Lin2020] it was expressed as a hope of the author that bootstrapping with positivity would eventually be used to find new critical points and phenomena of matrix models. In this paper, we do exactly that. As a proof of concept, we start off by using the bootstrap technique to estimate the critical point and exponent of the quartic single matrix model with known connections to 2d conformal field theory [di19952d].
@@
By exploiting the structure of the Schwinger-Dyson equations, we are able to give the first few terms of the free energy in the large $N$ precisely as a power series in the coupling constant at zero. For example, for the first two configurations we study, we find that
@@
which seems to have a search space of dimension two. We can begin to see evidence of critical behaviour similar to that of a cubic single matrix model. Just as in the 2-matrix models, we also compute the first few terms of a power series expansion for the free energy at $g=0$.
The article is organized as follows. In Section 2, we give the necessary background on matrix models, their critical behaviour, the Schwinger-Dyson equations, and bootstrapping with positivity. Next, in Section 3, we demonstrate how bootstrapping can be used to find the well-known critical point and exponent of the quartic single matrix model. In Section 4, we study the 2-matrix model mentioned above, and lastly in Section 5 we study the 3-matrix model. In Section 6, we summarize our work and its outlook. Examples of the Schwinger-Dyson equations and moments can be found for each model in the Appendices.
@@ 2.1 Matrix integrals
A matrix integral is an integral over a measurable set of matrices. In this paper, we restrict our interest to matrix integrals that are over some Cartesian power of the real vector space of $N\times N$ Hermitian matrices, denoted $\mathcal{H}_{N}$. There are two types of vastly different mathematical objects commonly referred to as a matrix integral or a matrix model.
First, there are convergent integrals over spaces of matrices, usually assigned an exponentially decaying weighted Lebesgue or Haar measure. One can use such matrix integrals to define probability distributions called matrix ensembles. Generally, (Hermitian) multi-matrix integrals are matrix integrals over some number of copies of $\mathcal{H}_{N}$, of the form
@@
where $S$ is some function such that the integral converges and each $dH_{i}$ is the Lebesgue measure on $\mathcal{H}_{N}$. With $\mathcal{Z}$ being a finite real number, we can define an associated matrix ensemble, sometimes called the Gibbs measure.
$\frac{1}{\mathcal{Z}}e^{-NS(H_{1},...,H_{m})}\prod_{i=1}^{m}dH_{i}$
In addition to computing the partition function $\mathcal{Z}$, one is often interested in computing the moments of such measures (see equations (2) and (3)) which can then be used to recover the distribution of eigenvalues [pastur2011eigenvalue]. Most results concerning such ensembles are found after taking the matrix size to go to infinity, with very few results for finite $N$. Such integrals have been studied extensively in the one matrix case for mostly polynomial potentials [ercolani2003asymptotics, johansson1998fluctuations, jonsson1998spectral], with applications to orthogonal polynomials and Log-Gases [deift2000orthogonal, forrester2010log]. Explicit results for multi-matrix integrals are far less common, partly due to the fact that most techniques concerning single matrix ensembles rely on some invariance of the measure to make a dimensional reduction of the integral.
The second kind of matrix integral is called a formal matrix integral, and is not an integral at all. A formal matrix integral is a formal series constructed by expanding all non-Gaussian terms of an expression like (1) around some coupling constants and then interchanging the order of integration and summation. Originally studied in [brezin1978planar], such formal series have connections to string theory, conformal field theory, quantum gravity, and combinatorics [di19952d, Lando2004]. In particular, in [brezin1978planar], it was realized that such integrals were the generating functions of maps i.e. graphs embedded onto orientable surfaces, considered up to orientation-preserving graph homeomorphisms. Such interpretations are essentially Feynman diagrammic expansions of finite dimensional integrals in Quantum Field Theory. The coefficients of such formal series grow double factorially, and therefore such formal series have a radius of convergence of zero. However, under general conditions, one can also consider such integrals as a formal Laurent series in the matrix size $N$. We say a matrix model has a genus expansion if for any word $W$ in the alphabet of matrix variables we can write
@@
Despite being fundamentally different objects, both convergent and formal matrix integrals have infinite systems of recursive relations they satisfy called the Schwinger-Dyson equations (SDE). If a matrix ensemble has a well-defined formal counterpart, often one finds that both sets of moments satisfy the same SDE in the large $N$ limit, which under general conditions can be shown to have a unique solution [guionnet2005combinatorial]. We will see that in this paper, the matrix models studied seem to each have a unique solution to the SDE that satisfies positivity constraints on their moments. It is worth noting here that the positivity constraints come from the Hamburger moment problem whose solutions form a convex set. Hence, if there is such a solution, it is either unique or there are infinitely many [schmudgen2020ten].
In this paper, we will exclusively study solutions of the SDE in the large $N$ limit. We denote mixed moments in the large $N$ limit as
$m_{W}:=\lim_{N\rightarrow\infty}\frac{1}{N}\langle\operatorname{Tr}W\rangle,$
and if one is only considering tracial moments from one matrix variable, we will denote them as
@@ 2.2 The Schwinger-Dyson equations
Let $W$ be a $m$-variate non-commutative polynomial in $\{H_{1},...,H_{m}\}$. Via Stoke’s formula, the following equality holds
@@
The SDE usually rely only on a finite set of moments and mixed moments to be solved, which can usually be deduced by utilizing a structural property of the SDE. In the above example, one needs precisely $m_{1},m_{2},...,m_{d-2}$ to solve for all other moments. In general this set of initial conditions is much harder to find, which brings us to the following notion. For a given range of the coupling constants, the search space of a matrix model is the minimum number of moments required as initial conditions for the model’s SDE. In order for this concept to be well-defined, it is of course required that such a model has a solution. As we will see, there are ranges of coupling constants in the models we study such that the SDE have no solution. At the time of writing this article there are no results establishing the existence or size of such search space multi-matrix models in any generality. General conditions required for SDE to have solutions were studied in [eynard2019solutions, guionnet2019asymptotics]. It is also very possible for a system of SDE to have more than one solution. Conditions for the existence of a unique solution are discussed in [guionnet2005combinatorial]. For more details on the derivation of the Schwinger-Dyson equations we refer the reader to [eynard2016counting, guionnet2019asymptotics].
@@ 2.3 Bootstrapping with positivity
The positivity constraints that can be derived for moments and mixed moments originates from the positivity of the spectral measure. This is easiest to see by starting with the Hamburger moment problem, which goes as follows: given a sequence of candidate real moments $(m_{0},m_{1},m_{2},...)$, does there exist a positive Borel measure $\mu$ on the real line whose moments correspond precisely to this sequence i.e.
@@
One can prove that such a probability measure exists if and only if the Hankel matrix of moments is positive definite:
@@
Bootstrapping with positivity refers to a process by which one combines the positivity constraints from the Hamburger moment problem with the relations given by the Schwinger-Dyson equations to derive explicit or numerical bounds on the moments of a matrix model. If the search space dimension of a model is low, one can derive simple expressions for the moments by solving the SDE up to some cutoff order, and then plugging them directly into the positivity constraints. The models studied in this paper all seem to have a search space dimension of one or two, making them ideal bootstrap candidates.
To derive positvity constraints for multi-matrix models, we must generalize these ideas. Consider a sequence of real numbers $\{m_{w}\}_{w\in\mathcal{A}}$ indexed by words $W$ in the alphabet formed from $\{H_{1},H_{2},...,H_{m}\}$. The sequence is called tracial if for any two cyclically equivalent words $w_{1}$ and $w_{2}$, we have that $m_{w_{1}}=m_{w_{2}}$. The necessary (but not sufficient) condition that a tracial sequence corresponds to mixed moments of a multi-matrix model is that the symmetric Hankel matrix $(m_{w_{1}^{*}w_{2}})_{w_{1},w_{2}}$ is positive semi-definite, i.e.
@@
Consider, for example, a 2-matrix model with matrix variables $A$ and $B$. All observables can be constructed in the real vector space spanned by the basis of lexigraphically ordered words in $A$ and $B$. This gives us the tracial sequence $\{1,m_{A},m_{B},m_{A^{2}},m_{AB},m_{B^{2}},\ldots\}$. Then, the following Hankel matrix is positive semi-definite:
$\left[\begin{array}[]{ccccccc}1&m_{A}&m_{B}&m_{A^{2}}&m_{AB}&m_{B^{2}}&\cdots\\ m_{A}&m_{A^{2}}&m_{AB}&m_{A^{3}}&m_{A^{2}B}&m_{AB^{2}}&\cdots\\ m_{B}&m_{BA}&m_{B^{2}}&m_{BA^{2}}&m_{BAB}&m_{B^{3}}&\cdots\\ m_{A^{2}}&m_{A^{3}}&m_{A^{2}B}&m_{A^{4}}&m_{A^{3}B}&m_{A^{2}B^{2}}&\cdots\\ m_{AB}&m_{BA^{2}}&m_{BAB}&m_{BA^{3}}&m_{BA^{2}B}&m_{BAB^{2}}&\cdots\\ m_{B^{2}}&m_{B^{2}A}&m_{B^{3}}&m_{B^{2}A^{2}}&m_{B^{2}AB}&m_{B^{4}}&\cdots\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\ddots\end{array}\right]\geq 0.$ (6)

### v2

# library.bib

@@ Abstract
Given a matrix model, by combining the Schwinger-Dyson equations with positivity constraints on its solutions, in the large $N$ limit one is able to obtain explicit and numerical bounds on its moments. This technique is known as bootstrapping with positivity. In this paper we use this technique to estimate the critical points and exponents of several multi-matrix models. As a proof of concept, we first show it can be used to find the well-studied quartic single matrix model’s critical phenomena. We then apply the method to several similar “unsolved” 2-matrix models with various quartic interactions. We conjecture and present strong evidence for the string susceptibility exponent for some of these models to be $\gamma=1/2$, which heuristically indicates that the continuum limit will likely be the Continuum Random Tree. For the other 2-matrix models, we find estimates of new string susceptibility exponents that may indicate a new continuum limit. We then study an unsolved 3-matrix model that generalizes the 3-colour model with cubic interactions. Additionally, for all of these models, we are able to derive explicitly the first several terms of the free energy in the large $N$ limit as a power series expansion in the coupling constants at zero by exploiting the structure of the Schwinger-Dyson equations.
@@ 1 Introduction
The primary quantities of interest in this paper are:
@@
Typically, one says that a model is solved (to leading order) if all these quantities can be written in terms of the coupling constants in the large $N$ limit. In general, these models are notoriously difficult to solve, even to leading order. For single matrix models with polynomial potentials and a small class of multi-matrix potentials, subleading terms can be obtained recursively using a process known as Topological Recursion [eynard2007invariants, eynard2016counting]. For general multi-matrix models, very little is known and much of it lacks a rigorous basis, with the exceptions of works like [guionnet2004first, guionnet2005combinatorial]. This is largely due to a shortage of analytic techniques. For an overview of what is known, see [kazakov2000solvable, eynard2011formal].
Of further interest is the study of the critical phenomena and a potential continuum limit of these models. Often, the quantities of interest have an asymptotic expansion around a critical point, i.e., a configuration of the coupling constants where one of the above quantities fails to be analytic. As a formal series that enumerates maps, such an expansion can be understood in terms of the asymptotic growth of the number of faces in the maps associated with the model. Such asymptotic expansions were first discovered in the 90’s and are related to 2D conformal field theory [di19952d, witten1990two]. This connection comes from studying the statistical properties of planar maps as random metric spaces. Such spaces are viewed as piecewise flat geometries used to construct models of 2d Euclidean quantum gravity. One then hopes to take appropriate limits so that the size of each polygon tends to zero and that these random metric spaces converge in some sense to a continuum theory. In the past twenty years, much work has been done to rigorously study such continuum limits [angel2003growth, chassaing2004random, le2010scaling]. Analogous ideas are used with Random Tensors [bonzom2012random, lionni2018colored] and Random Dynamical Triangulations [agishtein1992critical, loll2019quantum] for constructing toy models of higher dimensional gravity. It is worth noting that, despite matrix models corresponding to two-dimensional surfaces, multi-matrix models appear in both these frameworks as models with a space-time dimension greater than two [ambjorn2001lorentzian, lionni2018colored].
@@
where $f$ is a polynomial function and $2-\gamma\not\in\mathbb{Z}.$ This expansion is often referred to as a double scaling limit. The constant $\gamma$ is known as the string susceptibility exponent [di19952d]. When $\gamma=-1/2$, the associated continuum limit is expected to be the Brownian map [le2013uniqueness, marckert2006limit]. When $\gamma=1/2$, the associated continuum limit is expected to be the Continuum Random Tree [aldous1991continuumI, aldous1991continuumII, aldous1993continuumIII] which is known as branched polymers in the physics literature [cates1985fractal]. While these two are the most common and well-studied, other exponents are possible [kazakov1989appearance, korchemsky1992matrix]. In particular, this is possible when the model studied has multiple coupling constants that can be fine-tuned [distler19902d, ambjorn2016generalized]. Note that critical exponents do not always imply a continuum limit, but rather give a heuristically strong indication of them. For an example, see [kazakov1986ising].
To combat the above lack of analytical methods to solve multi-matrix models, the method of bootstrapping with positivity was first introduced in the innovative work [Lin2020]. Through this method, one derives successively stricter bounds on the moments of matrix models by combining the relations of the Schwinger-Dyson equations with positivity constraints on the spectral measure. Solving for these bounds is a non-linear optimization problem. In particular, in [kazakov2022analytic, zheng2023bootstrap] it was shown to be applicable to multi-matrix models that are unsolved, providing much-needed insight. In the same works, the relaxation bootstrap method was introduced, turning this non-linear problem into a linear one. Even more recently in [li2024analytic], further progress has been made in studying the 2-matrix model from [kazakov2022analytic] by adding an ansatz for the structure of moments. In [perez2024loop], positivity constraints on matrix integrals over the unitary group were used for bootstrapping. Bootstrapping with positivity has also been used in the previous work of the first two authors to study matrix models motivated by Noncommutative Geometry [hessam2022bootstrapping], ultimately leading to the recent analytic solution for many moments and the free energy of these models [khalkhali2024coloured]. The idea of bootstrapping with positivity constraints has also been used in other areas of mathematical physics such as matrix quantum mechanics [aikawa2022bootstrap, berenstein2022bootstrapping, berenstein2023semidefinite, bhattacharya2021numerical, tchoumakov2021bootstrapping], lattice gauge theory [kazakov2024bootstrap, kazakov2023bootstrap], Feynman integrals [zeng2023feynman], the Ising model on the lattice [cho2022bootstrapping], and more.
@@ 1.1 Outline of main results
In this paper, we will apply the bootstrapping with positivity technique to study the critical points and exponents of several multi-matrix models. In [Lin2020], it was expressed as a hope of the author that bootstrapping with positivity would eventually be used to find new critical points and phenomena in multi-matrix models. In this paper, we do exactly that. As a proof of concept, we start off by using the bootstrap technique to estimate the critical point and exponent of the quartic single matrix model with known connections to 2d conformal field theory [di19952d].
@@
By exploiting the structure of the Schwinger-Dyson equations, we are able to give the first few terms of the free energy in the large $N$ limit as a power series around zero of the coupling constant. For example, for the first two configurations we study, we find that
@@
In this model we begin to see evidence of critical behaviour similar to that of a cubic single matrix model. Just as in the 2-matrix models, we also compute the first few terms of a power series expansion for the free energy at $g=0$.
This article is organized as follows. In Section 2, we give the necessary background on matrix models, their critical behaviour, the Schwinger-Dyson equations, and bootstrapping with positivity. Next, in Section 3, we demonstrate how bootstrapping can be used to find the well-known critical point and exponent of the quartic single matrix model. In Section 4, we study the 2-matrix models mentioned above, and lastly in Section 5 we bootstrap the 3-matrix model. In Section 6, we summarize our work and its outlook. Examples of the Schwinger-Dyson equations and moments can be found for each model in the Appendices. These collections of equations are illustrative and not necessarily the entire list of equations used in bootstrapping. They contain clear patterns that may lead to hints at general analytical solutions.
@@ 2.1 Matrix integrals
In this paper, we restrict our interest to matrix integrals that are over some Cartesian power of the real vector space of $N\times N$ Hermitian matrices, denoted $\mathcal{H}_{N}$. There are two types of vastly different mathematical objects commonly referred to as a matrix integral or a matrix model.
First, there are convergent integrals over spaces of matrices, usually defined by an exponentially decaying matrix function. One can use such matrix integrals to define probability distributions called matrix ensembles. Generally, (Hermitian) multi-matrix integrals are matrix integrals over some number of copies of $\mathcal{H}_{N}$, of the form
@@
where $S$ is some function such that the integral converges and each $dH_{i}$ is the Lebesgue measure on $\mathcal{H}_{N}$. With $\mathcal{Z}$ being a finite real number, we can define an associated matrix ensemble, sometimes called the Gibbs measure
$\frac{1}{\mathcal{Z}}e^{-NS(H_{1},...,H_{m})}\prod_{i=1}^{m}dH_{i}.$
In addition to computing the partition function $\mathcal{Z}$, one is often interested in computing the moments of such measures (see equations (2) and (3)) which can then be used to recover the distribution of eigenvalues [pastur2011eigenvalue]. Most results concerning such ensembles are found after taking the matrix size to go to infinity, with very few results for finite $N$. Such integrals have been studied extensively in the one matrix case for mostly polynomial potentials [johansson1998fluctuations, jonsson1998spectral, ercolani2003asymptotics], with applications to orthogonal polynomials and Log-Gases [deift2000orthogonal, forrester2010log]. Explicit results for multi-matrix integrals are far less common, partly due to the fact that most techniques concerning single matrix ensembles rely on invariance of the measure to make a dimensional reduction of the integral.
The second kind of matrix integral is called a formal matrix integral which is a formal series constructed by expanding all non-Gaussian terms of an expression like (1) around some coupling constants and then interchanging the order of integration and summation. Originally studied in [brezin1978planar], such formal series have connections to string theory, conformal field theory, quantum gravity, and combinatorics [di19952d, Lando2004]. In particular, in [brezin1978planar], it was realized that such integrals were the generating functions of maps i.e. graphs embedded onto orientable surfaces, considered up to orientation-preserving graph homeomorphisms. Under general conditions, one can also consider such integrals as a formal Laurent series. We say a matrix model has a genus expansion if for any word $W$ in the alphabet of matrix variables we can write
@@
Despite being fundamentally different objects, both convergent and formal matrix integrals have infinite systems of recursive relations they satisfy called the Schwinger-Dyson equations (SDE). If a matrix ensemble has a well-defined formal counterpart, often one finds that both sets of moments satisfy the same SDE in the large $N$ limit, which under general conditions can be shown to have a unique solution [guionnet2005combinatorial]. The matrix models studied in this paper appear to each have a unique solution to the SDE that satisfies positivity constraints on their moments. It is worth noting here that the positivity constraints come from the Hamburger moment problem whose solutions form a convex set. Hence, if there is such a solution, it is either unique or there are infinitely many [schmudgen2020ten].
In this paper, we will exclusively study solutions of the SDE in the large $N$ limit. If one is only considering tracial moments from one matrix variable, we will denote them as
@@
Note that we can represent any word in two non-commuting matrices, as an integer sequence. For example, the word $AAAABBABBBB=A^{4}B^{2}AB^{4}$, can be represented as $(4,2,1,4)$. We denote mixed moments coming from a word $W$, by $m$ whose subscript is the associated sequence. For example the the moment for the word $AAAABBABBBB$ in th elarge $N$ limit is denoted $m_{4,2,1,4}$.
@@ 2.2 The Schwinger-Dyson equations
Let $W$ be a $m$-variate non-commutative polynomial in $\{H_{1},...,H_{m}\}$. Via Stokes’ formula, the following equality holds
@@
The SDE usually rely only on a finite set of moments and mixed moments to be solved, which can usually be deduced by utilizing a structural property of the SDE. In the above example, one needs precisely $m_{1},m_{2},...,m_{d-2}$ to solve for all other moments. In general this set of initial conditions is much harder to find, which brings us to the following notion. For a given range of the coupling constants, the search space of a matrix model is the minimum number of moments required as initial conditions for the model’s SDE. In order for this concept to be well-defined, it is of course required that such a model has a solution. As we will see, there are ranges of coupling constants in the models we study such that the SDE have no solution. At the time of writing this article there are no results establishing the existence or size of the search space of multi-matrix models in any generality. General conditions required for SDE to have solutions were studied in [eynard2019solutions, guionnet2019asymptotics]. It is also very possible for a system of SDE to have more than one solution. Conditions for the existence of a unique solution are discussed in [guionnet2005combinatorial]. For more details on the derivation of the Schwinger-Dyson equations we refer the reader to [eynard2016counting, guionnet2019asymptotics].
@@ 2.3 Bootstrapping with positivity
The positivity constraints that can be derived for moments and mixed moments originate from the positivity of the spectral measure. This is easiest to see by starting with the Hamburger moment problem, which goes as follows: given a sequence of candidate real moments $(m_{0},m_{1},m_{2},...)$, does there exist a positive Borel measure $\mu$ on the real line whose moments correspond precisely to this sequence i.e.
@@
One can prove that such a probability measure exists if and only if the Hankel matrix of moments is positive semi-definite:
@@
Bootstrapping with positivity refers to a process by which one combines the positivity constraints from the Hamburger moment problem with the relations given by the Schwinger-Dyson equations to derive explicit or numerical bounds on the moments of a matrix model. If the dimension of the search space of a model is small, one can derive simple expressions for the moments by solving the SDE up to some cutoff order, and then plugging them directly into the positivity constraints. The models studied in this paper all seem to have a search space dimension of one or two, making them ideal bootstrap candidates.
To derive positivity constraints for multi-matrix models, we must generalize these ideas. Consider a sequence of real numbers $\{m_{w}\}_{w\in\mathcal{A}}$ indexed by words $W$ in the alphabet formed from $\{H_{1},H_{2},...,H_{m}\}$. The sequence is called tracial if for any two cyclically equivalent words $w_{1}$ and $w_{2}$, we have that $m_{w_{1}}=m_{w_{2}}$. The necessary (but not sufficient) condition that a tracial sequence corresponds to mixed moments of a multi-matrix model is that the symmetric Hankel matrix $(m_{w_{1}^{*}w_{2}})_{w_{1},w_{2}}$ is positive semi-definite, i.e.
@@
Consider, for example, a 2-matrix model with matrix variables $A$ and $B$. All observables can be constructed in the real vector space spanned by the basis of lexicographically ordered words in $A$ and $B$. This gives us the tracial sequence $\{1,m_{1},m_{1},m_{2},m_{1,1},m_{2},\ldots\}$. Then, the following Hankel matrix is positive semi-definite:
$\left[\begin{array}[]{ccccccc}1&m_{1}&m_{1}&m_{2}&m_{1,1}&m_{2}&\cdots\\ m_{1}&m_{2}&m_{1,1}&m_{3}&m_{2,1}&m_{1,2}&\cdots\\ m_{1}&m_{1,1}&m_{2}&m_{1,2}&m_{1,1,1}&m_{3}&\cdots\\ m_{2}&m_{3}&m_{2,1}&m_{4}&m_{3,1}&m_{2,2}&\cdots\\ m_{1,1}&m_{1,2}&m_{1,1,1}&m_{1,3}&m_{1,21}&m_{1,1,2}&\cdots\\ m_{2}&m_{2,1}&m_{3}&m_{2,2}&m_{2,1,1}&m_{4}&\cdots\\ \vdots&\vdots&\vdots&\vdots&\vdots&\vdots&\ddots\end{array}\right]\geq 0.$ (5)